2.) 0.632 g of ammonium chloride (NH4Cl) are added to a 1 L solution of water at a temperature of 73°C in a tube separated by a semi-permeable membrane. Calculate the osmotic pressure of the solution.
1
Expert's answer
2020-05-14T14:33:26-0400

Solution:

Osmotic pressure is expressed by the formula:

Π = iMRT

where

Π is the osmotic pressure in atm

i = van 't Hoff factor of the solute

M = molar concentration in mol/L

R = universal gas constant = 0.08206 L·atm/mol·K

T = absolute temperature in K


T = 73°C + 273.15 = 346.15 K


Moles of NH4Cl = Mass of NH4Cl / Molar mass of NH4Cl

Molar mass of NH4Cl is 53.491 g/mol

Moles of NH4Cl = (0.632 g) / (53.491 g/mol) = 0.0118 mol


Molarity of NH4Cl = Moles of NH4Cl / Volume of solution

Molarity of NH4Cl = (0.0118 mol) / (1 L) = 0.0118 mol/L

The molar concentration of NH4Cl is 0.0118 mol/L.


Ammonium chloride dissociates completely in aqueous solution to produce ammonium cation and chloride anion:

NH4Cl → NH4+ + Cl-

Since salt (NH4Cl) dissociates into two ions, the value of the van’t Hoff factor here is 2.


Then,

Π = iMRT

Π = (2) × (0.0118 mol/L) × (0.08206 L·atm/mol·K) × (346.15 K) = 0.67 atm

Π = 0.67 atm


Answer: 0.67 atm is the osmotic pressure of the solution.